Please help me solve this question
The students of a school decided to beautify the school on Annual Day by fixing colourful flags on the straight passage of the school. They have 27 flags to be fixed at intervals if every 2m. The flags are stored at the position of the middle most flag. Ruchi was given the responsibility of placing the flags. Ruchi kept her books where the flags were stored. She could carry only one flag at a time. How much distance did she cover in completing this job and returning back to collect her books? What is the maximum distance she travelled carrying a flag?

Answer by rothauserc(4718)

The middle most flag has 13 flags to the right and 13 flags to the left, then
:
the total distance she travels is 2 * ( 2 + 4 + 6 + 8 + 10 + 12 + 14 + 16 + 18 + 20 + 22 + 24 + 26) = ?
:
the sum of the even integers is n(n+1), our n is 13, then
:
total distance = 2 * (13(13+1)) = 364
:
********************************************
Ruchi traveled a distance of 364m
********************************************
Ruchi traveled a maximum distance of 13 * 2 = 26m
*************************************************